# Stability for an inverse flux and an inverse boundary coefficient problems

### v1

@@ 1. Introduction and main results
Let $B\subset\mathbb{R}^{n}$ be a $C^{2}$ bounded open set with boundary $\mathcal{S}:=\partial B$ and set $U:=\mathbb{R}^{n}\setminus\overline{B}$. For convenience, we recall the following usual notations
@@
$\nu_{g}=(\nu_{g})^{j}\frac{\partial}{\partial x_{j}},\quad(\nu_{g})^{j}=\frac{g^{jk}\nu_{k}}{\sqrt{g^{\ell m}\nu_{\ell}\nu_{m}}}$
@@
Next, consider for $s\in\mathbb{R}$ and $r\in[1,\infty]$ the space
$B_{s,r}(\mathbb{R}^{n-1}):=\{w\in\mathscr{S}^{\prime};\;(1+|\xi|^{2})^{s/2}\widehat{w}\in L^{r}(\mathbb{R}^{n-1})\}$
@@
$\|w\|_{B_{s,r}(\mathbb{R}^{n-1})}:=\left\|(1+|\xi|^{2})^{s/2}\widehat{w}\right\|_{L^{r}(\mathbb{R}^{n-1})},$
where $\mathscr{S}^{\prime}$ denotes the space of tempered distributions on $\mathbb{R}^{n-1}$ and $\widehat{w}$ represents the Fourier transform of $w$. Utilizing local charts and a partition of unity, we construct $B_{s,r}(\mathcal{S})$ from $B_{s,r}(\mathbb{R}^{n-1})$. For further details, readers are referred to [6, p. 603] or [7, p. 3]. It follows from [6, Theorem 2.1] that the multiplication by $q\in B_{s,1}(\mathcal{S})$ defines a bounded operator on $H^{s}(\mathcal{S})$ with
(1.1) $\|qu\|_{H^{s}(\mathcal{S})}\leq\mathbf{c}_{0}\|q\|_{B_{s,1}(\mathcal{S})}\|u\|_{H^{s}(\mathcal{S})},$
where $\mathbf{c}_{0}:=2^{s/2}/(2\pi)^{n-1}$.
$B$$\mathcal{S}$$\Omega$$\Gamma$$U$
Figure 1. Illustration of the domain for Equation (1.2). $\mathcal{S}:=\partial B$, $B\Subset\Omega$, $\Gamma:=\partial\Omega$ and $U:=\mathbb{R}^{n}\setminus\overline{B}$.
Assume that $p\in L^{\infty}(U)$ and $q\in(B_{1/2,1}\cap L^{\infty})(\mathcal{S})$ such that $\Re q\geq 0$. Then we consider the exterior boundary-value problem
@@
Let $\Omega\Supset B$ be a $C^{0,1}$ bounded domain such that $\Omega\setminus\overline{B}$ is connected with boundary $\Gamma:=\partial\Omega$. We are interested in the inverse flux problem of determining the unknown function $\mathfrak{a}$ from the knowledge of $(u(0,\mathfrak{a})_{|\Gamma},\partial_{\nu_{g}}u(0,\mathfrak{a})_{|\Gamma})\in H^{3/2}(\Gamma)\times H^{1/2}(\Gamma)$, where $\nu_{g}$ denotes the outer unit normal to $\Gamma$ with respect to $g$. We refer to Figure 1 for an illustration of the domain related to the exterior boundary-value problem (1.2), as well as the associated measurement setup.
@@
[…]
@@
Let $40<\eta<1/$ 4. Then there exist constants $\mathbf{c}=\mathbf{c}(\zeta,\eta)>0$ and $c=c(\zeta,\eta)>0$ such that for all $\mathfrak{a}\in H^{1/2}(\mathcal{S})$ and $s\geq 1$ there holds
@@
Let $40<\eta<1/$ 4. There exist constants $\mathbf{c}=\mathbf{c}(\zeta,\eta)>0$ and $c=c(\zeta,\eta)>0$ such that for all $\mathfrak{a}\in H^{1/2}(\mathcal{S})\setminus\{0\}$ there holds the inequality
@@
where the norms $\|w\|_{L^{2}(\mathcal{S})}$ and $\|\nabla_{\tau_{g}}w\|_{L^{2}(\mathcal{S})}$ are defined similarly to $\|w\|_{L^{2}(\Gamma)}$.
@@
The same results are valid for the interior case as well. More specifically, let $B$ be a $C^{2}$ bounded domain and $\Omega\Subset B$ be a $C^{0,1}$ bounded open subset such that $B\setminus\overline{\Omega}$ is connected. When measurements are taken on $\Gamma:=\partial\Omega$, we can establish the same stability inequalities as those presented in Theorem 1.1 and Theorem 1.3. The interior problem of this nature has been addressed in Choulli-Takase [10].
@@ 2. Proof of the main results
Let $p\in L^{\infty}(U)$ and $q\in(B_{1/2,1}\cap L^{\infty})(\mathcal{S})$ satisfying $\Re p\geq\delta$ for some $\delta>0$ and $\Re q\geq 0$. If $(f,\mathfrak{a})\in L^{2}(U)\times L^{2}(\mathcal{S})$, then there exists a unique solution $u:=u(f,\mathfrak{a})\in H^{1}(U)$ to (1.2) and there holds the estimate
@@
If in addition to the above assumptions, $(g_{\ell m})\in W^{k+1,\infty}(\mathbb{R}^{n};\;\mathbb{R}^{n\times n})$, $\mathcal{S}\in C^{k+2}$, $p\in W^{k,\infty}(U)$, $q\in B_{k+1/2,1}(\mathcal{S})$ and $(f,\mathfrak{a})\in H^{k}(U)\times H^{k+1/2}(\mathcal{S})$ for $k\in\mathbb{N}\cup\{0\}$, then $u\in H^{k+2}(W\cap U)$ for any bounded open set $W\Supset B$ and we have
@@
[…]
@@
Next, we provide the proof of (2.2) in the specific case where $k=0$. The general case can be derived by induction with respect to $k$. For a bounded open set $W\Supset B$, we choose a $C^{1,1}$ bounded open set $W_{0}\Supset W$. Let $\chi\in C_{0}^{\infty}(W_{0})$ satisfying $\chi=1$ in a neighborhood of $W$. According to [13, Theorem 2.4.2.7], we have $\chi u\in H^{2}(W_{0}\cap U)$, where $u$ denotes the solution to (2.3). By applying [13, Theorem 2.3.3.6] to $\chi u$ in $W_{0}\cap U$ and utilizing (2.1), we obtain
@@
$\leq\mathbf{c}_{2}\left(\|f\|_{L^{2}(U)}+\|\mathfrak{a}\|_{H^{1/2}(\mathcal{S})}+\mathbf{c}_{0}\|q\|_{B_{1/2,1}(\mathcal{S})}\|u\|_{H^{1/2}(\mathcal{S})}\right)$
@@
Let $D:=\Omega\setminus\overline{B}$, $\mathfrak{a}\in H^{1/2}(\mathcal{S})$ and $u=u(0,\mathfrak{a})\in H^{2}(D)$. According to Theorem 1.1 in [11], for 3/$23/2<\eta<$ 2, there exist constants $\mathbf{c}=\mathbf{c}(g,p,B,\Omega,\eta)>0$ and $c=c(g,p,B,\Omega,\eta)>0$ such that for all $s\geq 1$, we have the inequality
@@ 3. Inverse boundary coefficient problem
In this section, we delve into an application of the aforementioned Theorems 1.1 and 1.3 in the context of an inverse boundary coefficient problem. Throughout the subsequent discussion, it is assumed that all functions are real-valued.
$B$$\mathcal{S}$$D=\Omega\setminus\overline{B}$$\Gamma$
The geometric setup for the problem we intend to investigate is rooted in the description provided in (1.2), but it is considered as an interior boundary value problem. For clarity, we recall the relevant details below. Let $n\geq 2$ be an integer, $B\subset\mathbb{R}^{n}$ be a $C^{2}$ bounded open set and $\Omega\Supset B$ be a $C^{2}$ bounded domain such that $D:=\Omega\setminus\overline{B}$ is connected. Set $\mathcal{S}:=\partial B$, $\Gamma:=\partial\Omega$ so that $\partial D=\mathcal{S}\cup\Gamma$ and define the set
$\mathscr{Q}_{k}:=\{q\in(B_{k+1/2,1}\cap L^{\infty})(\partial D);\;q\geq 0\;\text{and}\;q\not\equiv 0\},\quad k\in\mathbb{N}\cup\{0\}.$
@@
[…]
(see, for instance, [7, Lemma 2.9] for a similar approach), we can conclude that the boundary-value problem (3.1) admits a unique solution $u:=u_{q}(f,\mathfrak{a})\in H^{2}(D)$.
@@
$\mathscr{Q}:=\{q\in\mathscr{Q}_{k}\cap H^{1}(\partial D);\;\|q\|_{B_{k+1/2,1}(\partial D)}+\|q\|_{L^{\infty}(\partial D)}\leq\kappa\}.$
@@
There are numerous stability results associated to this inverse problem. For logarithmic stability estimates, encompassing single, double, and triple logarithms, readers are directed to the following references: [1, 3, 5, 8, 16, 9, 7]. For stronger stability estimates, Choulli [6, Theorem 1.2] establishes a Hölder type stability estimate when observation data are collected on the same boundary as the unknown function. Additionally, Hu and Yamamoto [14] demonstrate a Hölder type stability estimate in the two-dimensional case when the unknown functions are constrained to be analytic or piecewise constant. Chaabane and Jaoua [4] establish a Lipschitz type stability estimate under a monotony assumption concerning the unknown coefficients, while Sincich [15] proves another Lipschitz type stability estimate when the unknown coefficients are assumed to be piecewise constant functions. Furthermore, Yang, Choulli, and Cheng [17] propose a numerical scheme employing the boundary-element method for the specific scenario where $D$ is an annulus.
@@
Let $k>n/2$ be an integer and $K>0$. Assume that $B$ and $\Omega$ are of class $C^{k+2}$ and $(g_{\ell m})\in W^{k+1,\infty}(\mathbb{R}^{n};\mathbb{R}^{n\times n})$. Fix $(f,\mathfrak{a})\in H^{k}(D)\times H^{k+1/2}(\partial D)\setminus\{(0,0)\}$ such that $f\geq 0$ and $\mathfrak{a}\geq 0$. Then, there exists a constant $\mathbf{c}=\mathbf{c}(\varsigma,K)>0$ such that for all $q_{j}\in\mathscr{Q}$ and $u_{j}=u_{q_{j}}(f,\mathfrak{a})\in H^{k+2}(D)$ satisfying (3.1) for $j=1,2$, there holds
@@
Before proving Theorem 3.1, we present some of its consequences. Let $(\lambda_{m})$ be the sequence of eigenvalues of the Laplace-Beltrami operator $-\Delta_{\tau_{g}}$ on $\mathcal{S}$, satisfying
$0=\lambda_{0}<\lambda_{1}\leq\lambda_{2}\leq\cdots\quad\text{and}\quad\lim_{m\to\infty}\lambda_{m}=\infty.$
@@
$\|\nabla_{\tau_{g}}(q_{1}-q_{2})\|_{L^{2}(\mathcal{S})}=\sum_{\lambda_{m}\leq\lambda}\lambda_{m}|a_{m}|^{2}\leq\lambda\|q_{1}-q_{2}\|_{L^{2}(\mathcal{S})}.$
@@
Let $k>n/2$ be an integer. Assume that $B$ and $\Omega$ are of class $C^{k+2}$ and that $(g_{\ell m})\in W^{k+1,\infty}(\mathbb{R}^{n};\mathbb{R}^{n\times n})$. Fix $(f,\mathfrak{a})\in H^{k}(D)\times H^{k+1/2}(\partial D)\setminus\{(0,0)\}$ such that $f\geq 0$ and $\mathfrak{a}\geq 0$. Then, there exist constants $\mathfrak{c}_{1}=\mathfrak{c}_{1}(\varsigma)>0$ and $\mathfrak{c}_{2}=\mathfrak{c}_{2}(\varsigma)>0$ such that for all $q\in\mathscr{Q}$ and $u_{q}=u_{q}(f,\mathfrak{a})\in H^{k+2}(D)$ satisfying (3.1), there hold
@@
Let $u_{\kappa}:=u_{\kappa}(f,\mathfrak{a})\in C^{2}(\overline{D})$ be the unique solution of (3.1) with $q:=\kappa$. We will prove that $\min_{\overline{D}}u_{\kappa}>0$. In the case where $u_{\kappa}$ is equal to a constant, it is nonzero by $(f,\mathfrak{a})\neq(0,0)$. Therefore, from $\kappa u_{\kappa}=\mathfrak{a}\geq 0$ on $\partial D$, we deduce that $\mathfrak{a}$ and $u_{\kappa}=\mathfrak{a}/\kappa>0$ are both equal to positive constants. When $u_{\kappa}$ is not equal to a constant, $u_{\kappa}$ reaches its minimum at some $x_{\kappa}\in\partial D$ according to the strong maximum principle (e.g., [12, Theorem 3.5]). We have $u_{\kappa}(x_{\kappa})>0$ because otherwise we will have
@@
which contradicts $\partial_{\nu_{g}}u_{\kappa}(x_{\kappa})<0$ by [2, Lemma 3.1].
@@
Since $(\kappa-q)u_{\kappa}\geq\left(\kappa-\|q\|_{L^{\infty}(\partial D)}\right)u_{\kappa}(x_{\kappa})\geq 0$ on $\partial D$, proceeding as above, we obtain $v\geq 0$ in $\overline{D}$. Thus, $\min_{\overline{D}}u_{q}\geq\min_{\overline{D}}u_{\kappa}>0$. which means that (3.4) holds with $\mathfrak{c}_{2}:=\min_{\overline{D}}u_{\kappa}>0$. ∎
@@
Let $k>n/2$ be an integer and $40<\eta<1/$ 4. Assume that $B$ and $\Omega$ are of class $C^{k+2}$ and that $(g_{\ell m})\in W^{k+1,\infty}(\mathbb{R}^{n};\mathbb{R}^{n\times n})$. Fix $(f,\mathfrak{a})\in H^{k}(D)\times H^{k+1/2}(\partial D)\setminus\{(0,0)\}$ such that $f\geq 0$ and $\mathfrak{a}\geq 0$. Then there exist constants $\mathbf{c}=\mathbf{c}(\varsigma,\eta)>0$ and $c=c(g,\kappa,B,\Omega,\eta)>0$ such that for all $q_{j}\in\mathscr{Q}$ with $u_{j}=u_{q_{j}}(f,\mathfrak{a})\in H^{k+2}(D)$ satisfying (3.1) for $j=1,2$, and $s\geq 1$ there holds

### v2

@@ 1. Introduction and main results
Let $B\subset\mathbb{R}^{n}$ be a smooth bounded open set with boundary $\mathcal{S}:=\partial B$ and set $U:=\mathbb{R}^{n}\setminus\overline{B}$. For convenience, we recall the following usual notations
@@
$\nu_{g}=(\nu_{g})^{j}\frac{\partial}{\partial x_{j}},\quad(\nu_{g})^{j}=\frac{g^{jk}\nu_{k}}{\sqrt{g^{\ell m}\nu_{\ell}\nu_{m}}},$
@@
Next, consider for $t\in(0,\infty)\setminus\mathbb{N}$ the space
$H^{t}(\mathbb{R}^{n-1}):=\{w\in\mathscr{S}^{\prime};\;(1+|\xi|^{2})^{t/2}\widehat{w}\in L^{2}(\mathbb{R}^{n-1})\}$
@@
$\|w\|_{H^{t}(\mathbb{R}^{n-1})}:=\left\|(1+|\xi|^{2})^{t/2}\widehat{w}\right\|_{L^{2}(\mathbb{R}^{n-1})},$
where $\mathscr{S}^{\prime}$ denotes the space of tempered distributions on $\mathbb{R}^{n-1}$ and $\widehat{w}$ represents the Fourier transform of $w$. Utilizing local charts and a partition of unity, we construct $H^{t}(\mathcal{S})$ from $H^{t}(\mathbb{R}^{n-1})$. For further details, readers are referred to [Agranovich2015, Section 2.2]. It follows from [Agranovich2015, Theorem 1.9.2 and Theorem 2.3.6] that the multiplication by $q\in C^{\lfloor t\rfloor,1}(\mathcal{S})$, where $\lfloor t\rfloor$ denotes the largest integer not exceeding $t$, defines a bounded operator on $H^{t}(\mathcal{S})$ with
(1.1) $\|qu\|_{H^{t}(\mathcal{S})}\leq\mathbf{c}_{0}\|q\|_{C^{\lfloor t\rfloor,1}(\mathcal{S})}\|u\|_{H^{t}(\mathcal{S})},$
where $\mathbf{c}_{0}=\mathbf{c}_{0}(n,t,\mathcal{S})>0$ is a constant.
Assume that $p\in L^{\infty}(U)$ and $q\in C^{0,1}(\mathcal{S})$ such that $\Re q\geq 0$. Then we consider the exterior boundary-value problem
@@
Let $\Omega\Supset B$ be a $C^{0,1}$ bounded domain such that $\Omega\setminus\overline{B}$ is connected with boundary $\Gamma:=\partial\Omega$. We consider (1.2) for the case $f=0$ and are interested in the inverse flux problem of determining the unknown function $\mathfrak{a}$ from the knowledge of $(u(0,\mathfrak{a})_{|\Gamma},\partial_{\nu_{g}}u(0,\mathfrak{a})_{|\Gamma})\in H^{3/2}(\Gamma)\times H^{1/2}(\Gamma)$, where $\nu_{g}$ denotes the outer unit normal to $\Gamma$ with respect to $g$. We refer to Figure 1 for an illustration of the domain related to the exterior boundary-value problem (1.2), as well as the associated measurement setup.
[…]
Figure 1. Illustration of the domain for Equation (1.2). $\mathcal{S}:=\partial B$, $B\Subset\Omega$, $\Gamma:=\partial\Omega$ and $U:=\mathbb{R}^{n}\setminus\overline{B}$. $\Gamma$ is an accessible and $\mathcal{S}$ is an inaccessible part of the boundary.
@@
$\|w\|_{L^{2}(\Gamma)}$ $:=\left(\int_{\Gamma}|w|^{2}\,dS_{g}\right)^{1/2},$
$\|w\|_{H^{1}(\Gamma)}$ $:=\left(\int_{\Gamma}|w|^{2}\,dS_{g}+\int_{\Gamma}|\nabla_{\tau_{g}}w|^{2}\,dS_{g}\right)^{1/2}.$
@@
Let $0<\eta<1/4$. Then there exist constants $\mathbf{c}=\mathbf{c}(\zeta,\eta)>0$ and $c=c(\zeta,\eta)>0$ such that for all $\mathfrak{a}\in H^{1/2}(\mathcal{S})$ and $s\geq 1$ there holds
@@
Let $0<\eta<1/4$. There exist constants $\mathbf{c}=\mathbf{c}(\zeta,\eta)>0$ and $c=c(\zeta,\eta)>0$ such that for all $\mathfrak{a}\in H^{1/2}(\mathcal{S})\setminus\{0\}$ there holds the inequality
@@
where the norms $\|w\|_{L^{2}(\mathcal{S})}$ and $\|\nabla_{\tau_{g}}w\|_{L^{2}(\mathcal{S})}$ are defined similarly to $\|w\|_{L^{2}(\Gamma)}$. Roughly speaking, $\mathcal{A}$ imposes a constraint on the high-frequency components. Indeed, as shown in Corollary 3.2, finite-dimensional subspaces are contained in $\mathcal{A}$.
@@
The same results are valid for the interior case as well. More specifically, let $B$ be a smooth bounded domain and $\Omega\Subset B$ be a $C^{0,1}$ bounded open subset such that $B\setminus\overline{\Omega}$ is connected. When measurements are taken on $\Gamma:=\partial\Omega$, we can establish the same stability inequalities as those presented in Theorem 1.1 and Theorem 1.3. The interior problem of this nature has been addressed in Choulli-Takase [Choulli2025].
@@ 2. Proof of the main results
Let $p\in L^{\infty}(U)$ and $q\in C^{0,1}(\mathcal{S})$ satisfying $\Re p\geq\delta$ for some $\delta>0$ and $\Re q\geq 0$. If $(f,\mathfrak{a})\in L^{2}(U)\times L^{2}(\mathcal{S})$, then there exists a unique solution $u:=u(f,\mathfrak{a})\in H^{1}(U)$ to (1.2) and there holds the estimate
@@
If in addition to the above assumptions, $(g_{\ell m})\in W^{k+1,\infty}(\mathbb{R}^{n};\mathbb{R}^{n\times n})$, $p\in W^{k,\infty}(U)$, $q\in C^{k,1}(\mathcal{S})$ and $(f,\mathfrak{a})\in H^{k}(U)\times H^{k+1/2}(\mathcal{S})$ for $k\in\mathbb{N}\cup\{0\}$, then $u\in H^{k+2}(W\cap U)$ for any bounded open set $W\Supset B$ and we have
@@
$h(u,v):=\int_{U}(\langle\nabla_{g}u,\nabla_{g}\overline{v}\rangle+pu\overline{v})dV_{g}+\int_{\mathcal{S}}qu\overline{v}dS_{g},\quad u,v\in H^{1}(U)$
@@
Next, we provide the proof of (2.2) in the specific case where $k=0$. The general case can be derived by induction with respect to $k$. For a bounded open set $W\Supset B$, we choose a $C^{1,1}$ bounded open set $W_{0}\Supset W$. Let $\chi\in C_{0}^{\infty}(W_{0})$ satisfying $\chi=1$ in a neighborhood of $W$. According to [Grisvard1985, Theorem 2.4.2.7], we have $\chi u\in H^{2}(W_{0}\cap U)$, where $u$ denotes the solution to (2.3). By applying [Grisvard1985, Theorem 2.3.3.6] to $\chi u$ in $W_{0}\cap U$ and utilizing (2.1), we obtain
@@
$\leq\mathbf{c}_{2}\left(\|f\|_{L^{2}(U)}+\|\mathfrak{a}\|_{H^{1/2}(\mathcal{S})}+\mathbf{c}_{0}\|q\|_{C^{0,1}(\mathcal{S})}\|u\|_{H^{1/2}(\mathcal{S})}\right)$
@@
Let $D:=\Omega\setminus\overline{B}$, $\mathfrak{a}\in H^{1/2}(\mathcal{S})$ and $u=u(0,\mathfrak{a})\in H^{2}(D)$. According to [CT2024b, Theorem 1.1], for $3/2<\eta<2$, there exist constants $\mathbf{c}=\mathbf{c}(g,p,B,\Omega,\eta)>0$ and $c=c(g,p,B,\Omega,\eta)>0$ such that for all $s\geq 1$, we have the inequality
@@ 3. Inverse boundary coefficient problem
In this section, we delve into an application of the aforementioned Theorems 1.1 and 1.3 in the context of an inverse boundary coefficient problem. Throughout the subsequent discussion, it is assumed that all functions are real-valued for simplicity. The reason is that, while a parallel argument can be made for complex-valued functions by decomposing them into their real and imaginary parts, this leads to rather cumbersome notation.
The geometric setup for the problem we intend to investigate is rooted in the description provided in (1.2), but it is considered as an interior boundary value problem. For clarity, we recall the relevant details below. Let $n\geq 2$ be an integer, $B\subset\mathbb{R}^{n}$ be a smooth bounded open set and $\Omega\Supset B$ be a smooth bounded domain such that $D:=\Omega\setminus\overline{B}$ is connected. Set $\mathcal{S}:=\partial B$, $\Gamma:=\partial\Omega$ so that $\partial D=\mathcal{S}\cup\Gamma$ and define the set
$\mathscr{Q}_{k}:=\{q\in C^{k,1}(\partial D);\;q\geq 0\;\text{and}\;q\not\equiv 0\},\quad k\in\mathbb{N}\cup\{0\}.$
@@
$h(u,v):=\int_{D}\langle\nabla_{g}u,\nabla_{g}v\rangle dV_{g}+\int_{\partial D}quvdS_{g},\quad u,v\in H^{1}(D),$
(see, for instance, [Choulli2016, Lemma 2.9] for a similar approach), we can conclude that the boundary-value problem (3.1) admits a unique solution $u:=u_{q}(f,\mathfrak{a})\in H^{2}(D)$.
@@
$\mathscr{Q}:=\{q\in\mathscr{Q}_{k};\;\|q\|_{C^{k,1}(\partial D)}\leq\kappa\}.$
@@
There are numerous stability results associated to this inverse problem. For logarithmic stability estimates, encompassing single, double, and triple logarithms, readers are directed to the following references: [Alessandrini2003, Chaabane2004, Cheng2008, Choulli2018, Sincich2015, Choulli2016a, Choulli2016]. For stronger stability estimates, Choulli [Choulli2002, Theorem 1.2] establishes a Hölder type stability estimate when observation data are collected on the same boundary as the unknown function. Additionally, Hu and Yamamoto [Hu2015] demonstrate a Hölder type stability estimate in the two-dimensional case when the unknown functions are constrained to be analytic or piecewise constant. Chaabane and Jaoua [Chaabane1999] establish a Lipschitz type stability estimate under a monotony assumption concerning the unknown coefficients, while Sincich [Sincich2007] proves another Lipschitz type stability estimate when the unknown coefficients are assumed to be piecewise constant functions. Furthermore, Yang, Choulli, and Cheng [Yang2005] propose a numerical scheme employing the boundary-element method for the specific scenario where $D$ is an annulus.
@@
Let $k>n/2$ be an integer and $K>0$. Assume that $(g_{\ell m})\in W^{k+1,\infty}(\mathbb{R}^{n};\mathbb{R}^{n\times n})$. Fix $(f,\mathfrak{a})\in H^{k}(D)\times H^{k+1/2}(\partial D)\setminus\{(0,0)\}$ such that $f\geq 0$ and $\mathfrak{a}\geq 0$. Then, there exists a constant $\mathbf{c}=\mathbf{c}(\varsigma,K)>0$ such that for all $q_{j}\in\mathscr{Q}$ and $u_{j}=u_{q_{j}}(f,\mathfrak{a})\in H^{k+2}(D)$ satisfying (3.1) for $j=1,2$, there holds
@@
Before proving Theorem 3.1, we present some of its consequences. Let $(\lambda_{m})$ be the sequence of eigenvalues of the Laplace-Beltrami operator $\Delta_{\tau_{g}}$ on $\mathcal{S}$, satisfying
$0=\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\leq\cdots\quad\text{and}\quad\lim_{m\to\infty}\lambda_{m}=\infty.$
@@
$\|\nabla_{\tau_{g}}(q_{1}-q_{2})\|_{L^{2}(\mathcal{S})}=\sqrt{\sum_{\lambda_{m}\leq\lambda}\lambda_{m}|a_{m}|^{2}}\leq\sqrt{\lambda}\|q_{1}-q_{2}\|_{L^{2}(\mathcal{S})}.$
@@
Let $k>n/2$ be an integer. Assume that $(g_{\ell m})\in W^{k+1,\infty}(\mathbb{R}^{n};\mathbb{R}^{n\times n})$. Fix $(f,\mathfrak{a})\in H^{k}(D)\times H^{k+1/2}(\partial D)\setminus\{(0,0)\}$ such that $f\geq 0$ and $\mathfrak{a}\geq 0$. Then, there exist constants $\mathfrak{c}_{1}=\mathfrak{c}_{1}(\varsigma)>0$ and $\mathfrak{c}_{2}=\mathfrak{c}_{2}(\varsigma)>0$ such that for all $q\in\mathscr{Q}$ and $u_{q}=u_{q}(f,\mathfrak{a})\in H^{k+2}(D)$ satisfying (3.1), there hold
@@
Let $u_{\kappa}:=u_{\kappa}(f,\mathfrak{a})\in C^{2}(\overline{D})$ be the unique solution of (3.1) with $q:=\kappa$. We will prove that $\min_{\overline{D}}u_{\kappa}>0$. In the case where $u_{\kappa}$ is equal to a constant, it is nonzero by $(f,\mathfrak{a})\neq(0,0)$. Therefore, from $\kappa u_{\kappa}=\mathfrak{a}\geq 0$ on $\partial D$, we deduce that $\mathfrak{a}$ and $u_{\kappa}=\mathfrak{a}/\kappa>0$ are both equal to positive constants. When $u_{\kappa}$ is not equal to a constant, $u_{\kappa}$ reaches its minimum at some $x_{\kappa}\in\partial D$ according to the strong maximum principle (e.g., [Gilbarg2001, Theorem 3.5]). We have $u_{\kappa}(x_{\kappa})>0$ because otherwise we will have
@@
which contradicts $\partial_{\nu_{g}}u_{\kappa}(x_{\kappa})<0$ by [Ammari2020, Lemma 3.1].
@@
Since $(\kappa-q)u_{\kappa}\geq\left(\kappa-\|q\|_{C(\partial D)}\right)u_{\kappa}(x_{\kappa})\geq 0$ on $\partial D$, proceeding as above, we obtain $v\geq 0$ in $\overline{D}$. Thus, $\min_{\overline{D}}u_{q}\geq\min_{\overline{D}}u_{\kappa}>0$, which means that (3.4) holds with $\mathfrak{c}_{2}:=\min_{\overline{D}}u_{\kappa}>0$. ∎
@@
Let $k>n/2$ be an integer and $0<\eta<1/4$. Assume that $(g_{\ell m})\in W^{k+1,\infty}(\mathbb{R}^{n};\mathbb{R}^{n\times n})$. Fix $(f,\mathfrak{a})\in H^{k}(D)\times H^{k+1/2}(\partial D)\setminus\{(0,0)\}$ such that $f\geq 0$ and $\mathfrak{a}\geq 0$. Then there exist constants $\mathbf{c}=\mathbf{c}(\varsigma,\eta)>0$ and $c=c(g,\kappa,B,\Omega,\eta)>0$ such that for all $q_{j}\in\mathscr{Q}$ with $u_{j}=u_{q_{j}}(f,\mathfrak{a})\in H^{k+2}(D)$ satisfying (3.1) for $j=1,2$, and $s\geq 1$ there holds